We've seen for some time now that one of the PRM OSEM signals has been gone, and all of the SRM signals seem dark. We had tried squishing various cables to no avail.
Today I played some "musical satellite boxes," in an attempt to see if the problems are in the chambers or in the signal chains. That is, I swapped the OSEM cables from the vacuum feedthroughs between the satellite boxes, and observed what happened.
It seems clear that something is up with SRM inside the chamber. For PRM, it's not so clear...
Somehow, issues with the LR channel follow both the PRM OSEMs and the PRM satellite box.
PRM LR first went dark on Jul 2nd, after the IFO was vented, but before we took any doors off (which happened on the 5th). I'm not sure what may have caused this.
SRM OSEMS first went dark on the evening of Jul 18, the day before ELOG 12310, when ITMY was moved in the same chamber. Maybe this ELOG was written about work the day before, but the sensors show disturbances over the course of hours. I think we need to double check the connections in chamber.